1 An ion with mass number 56 contains 3 units of positive charge and 30.4% more neutrons than

electrons. Assign symbol to this ion.

e 2+He ®¾¾ He - + +Q.2 Calculate the energy required for the process:

Given, the ionization energy for H-atom in ground state is 2.18 x 10-18 Jatom-1

Grade:11

1 Answers

varun
7 years ago
Let the number of electrons in ion=X
no of protons or atomic number of element =X+3
no of neutrons in ion =X+30.4% of X
                              =X+0.304X
Atomic mass of element in sum of the element of neutrons and no of protons of the elememt,,thus 
56=(X+3)+(1.304)X
solving.X=3
hence atomic no of element is 23+3=26
threfore the element is IRON
BEST OF LUCK 
spare me if the answer is wrong